OpenStudy (anonymous):

Bob has two savings accounts. He deposited $100 more into account B than account A. After a period of time, account A has earned $105 in interest at 7%, and account B has earned $80 in interest at a rate of 5%. 1. Write an equation to represent the situation. Explain each variable used. 2. How much money did Bob initially deposit into each account? Solve the equation. I dont understand this..Please help!!

OpenStudy (anonymous):

Please help me! I need to know this before the test!

hero (hero):

B = A + 100 .07A + .05B = 185

OpenStudy (anonymous):

What do I do after that?

hero (hero):

Use substitution to substitute the expression for B into the second equation: .07A + .05(A + 100) = 185

OpenStudy (anonymous):

Oh I think I see...Do I subsitute?

hero (hero):

After substitution, you solve for A, then for B

OpenStudy (anonymous):

0.12A=85?

hero (hero):

I think you rushed it.

hero (hero):

You forgot to distribute the .05

OpenStudy (anonymous):

Oh Ok, now How do I get the B part?

hero (hero):

@MattzSkatinn...one step at a time. Try solving for A first.

OpenStudy (anonymous):

I think A=1500

hero (hero):

Okay, so according to the original formulas, B will be $100 more, right?

OpenStudy (anonymous):

Soo..1600?

hero (hero):

Good job bro

hero (hero):

By the way, the first formula represents the amount of B in terms of A The second formula represents total interest.

OpenStudy (anonymous):

Thanks so much! I think I understand it now!

OpenStudy (anonymous):

That I dont understand..

OpenStudy (anonymous):

So he deposited 1500 into A and 1600 into B? @Hero

hero (hero):

Yes

OpenStudy (anonymous):

Ok, and I dont understand what you meant by what the formulas represent

hero (hero):

A = amount invested in account A B = amount invested in Account B He deposited $100 more into account B than account A: B = A + 100 Account A has earned $105 in interest at 7%, and account B has earned $80 in interest at a rate of 5%: .07A + .05B = 185

hero (hero):

Maybe that will help you understand better. Each sentence can be translated into an equation.

hero (hero):

The first equation is obvious. The trick to the second equation is realizing that 105 + 80 gives you total interest. Once this is realized, you try to write an equation for total interest, or rather an equation that expresses 185 in terms of accounts A and B.

OpenStudy (anonymous):

Ok I think I understand..

hero (hero):

Can also be solved using I = PRT if you're familiar with that